OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

OpenStudy (anonymous):

Point-slope form: y - y1 = m(x - x1)

OpenStudy (anonymous):

Figure out the slope and plug it in as m. Then pick a point and plug that into y1 and x1.

OpenStudy (anonymous):

The slope of a line is the ratio of the change in y over the change in x. Slope/m = (y2 - y1)/ (x2 - x1)

OpenStudy (anonymous):

Plug the points you have into that and get slope.

OpenStudy (anonymous):

2 - 1 / - 3 - 2 = ?